# Beyond “Not Novel Enough”: Enriching Scholarly Critique with LLM-Assisted Feedback

### v3

@@ Abstract
Novelty assessment is a central yet understudied aspect of peer review, particularly in high-volume fields like NLP where reviewer capacity is increasingly strained. We present a structured approach for automated novelty evaluation that models expert reviewer behavior through three stages: content extraction from submissions, retrieval and synthesis of related work, and structured comparison for evidence-based assessment. Our method is informed by analysis of human-written novelty reviews and captures key patterns such as independent claim verification and contextual reasoning. Evaluated on 182 ICLR 2025 submissions with human annotated reviewer novelty assessments, the approach achieves 86.5% alignment with human reasoning and 75.3% agreement on novelty conclusions—substantially outperforming existing LLM-based baselines. The method produces detailed, literature-aware analyses and improves consistency over ad hoc reviewer judgments. These results highlight the potential for structured LLM-assisted approaches to support more rigorous and transparent peer review without displacing human expertise. Data and code are made available.  [^1] https://github.com/UKPLab/arxiv2025-assessing-paper-novelty/
@@ 1 Introduction
The peer review system is collapsing under its own success. Two independent committees at NeurIPS 2021 disagreed on 23% of identical papers beygelzimer2023machinelearningreviewprocess—a breakdown in consistency that signals deeper problems than mere capacity constraints. With manuscript submissions doubling every 15 years RePEc:spr:scient:v:84:y:2010:i:3:d:10.1007_s11192-010-0202-z and reviewers now handling 14 evaluations annually Diaz2024Annotations, the system’s 15 million annual reviewing hours Aczel2021 are producing increasingly unreliable outcomes.
@@
Among peer review tasks, novelty assessment stands out as one of the most problematic Ernst2020UnderstandingPR Horbach2018TheAO. Novelty assessment requires reviewers to determine whether a submission makes sufficiently original contributions by identifying what specific advances it makes beyond existing work, evaluating whether these advances are significant enough to warrant publication, and verifying that the authors have accurately characterized their contributions relative to prior research. This knowledge-intensive process demands that reviewers maintain comprehensive awareness of related work across their field and can precisely distinguish between meaningful innovations and incremental modifications—a task that becomes exponentially more difficult as publication rates accelerate and research domains specialize. Overwhelmed reviewers often resort to superficial analyses, producing vague feedback like "not novel enough" without clear justification. The challenge compounds when reviewers encounter papers outside their specific expertise, leading to either overly conservative rejections or inadequate assessments that fail to catch incremental work kuznetsov2024naturallanguageprocessingpeer.
@@
Recent advances in large language models present an unprecedented opportunity to address these novelty assessment challenges at scale. These breakthrough technologies have revolutionized text processing and demonstrated remarkable performance across knowledge-intensive tasks Raiaan2024ARO, with recent technical advancements expanding capabilities to specialized reasoning and efficient inference Li2024LLMIS; Zhang2025KAGThinkerIT.
@@
While recent LLM advances create this opportunity, no existing work specifically addresses novelty assessment as a dedicated task within the peer review process. Prior research incorporates novelty evaluation within idea generation pipelines radensky2025scideatorhumanllmscientificidea; lu2024aiscientist; li2024chainideasrevolutionizingresearch, generates peer reviews with novelty assessments occurring as a result of them existing in peer reviews from training data idahl-ahmadi-2025-openreviewer; darcy2024margmultiagentreviewgeneration, or adds novelty assessment steps to review synthesis pipelines for improvement zhu2025deepreviewimprovingllmbasedpaper. However, these approaches either operate on synthetic ideas rather than real research contributions or fail to evaluate novelty assessment capabilities in isolation. This represents a critical gap requiring specialized methodologies for peer review novelty assessment.
@@ 2 Related Work
Our work is positioned at the peer review stage of scientific research, where our system operates when a manuscript is submitted for evaluation. While previous works darcy2024margmultiagentreviewgeneration idahl-ahmadi-2025-openreviewer zhu2025deepreviewimprovingllmbasedpaper chitale2025autorev Chang2025TreeReviewAD nemecek2025feasibilitytopicbasedwatermarkingacademic have developed end-to-end peer review generation pipelines that may implicitly include novelty assessment steps, we are the first to focus specifically on building a dedicated pipeline for novelty assessment and the first to systematically evaluate LLMs on this task. A related line of work operates at the ideation stage of research radensky2025scideatorhumanllmscientificidea shahid2025literature li2024chainideasrevolutionizingresearch lu2024aiscientist, developing pipelines for research idea generation that aim to improve novelty through feedback loops from a novelty assessor. In contrast, we operate at a more mature stage where ideas have been fully executed and comparative analyses are well-formulated. The evaluation in ideation-stage works focuses on synthetic ideas that are typically abstract and loosely defined, whereas we evaluate concrete, polished research contributions that have undergone the refinement process of execution and manuscript preparation.
@@
Our work employs an extensive related work discovery pipeline that collects papers cited within the submission and additionally retrieves related papers by querying with prompts generated by GPT-4.1. Papers are then ranked using an embedding-based method and reranked using RankGPT. We adapt this general approach from existing work radensky2025scideatorhumanllmscientificideashahid2025literatureli2024chainideasrevolutionizingresearch with modifications to ranking and filtering for our specific task. Similar retrieval-rank-rerank pipelines have been used for related work generation agarwal2025litllmtoolkitscientificliterature. Another retrieval approach is OpenScholar asai2024openscholarsynthesizingscientificliterature, which uses an LLM-RAG based approach to answer scientific queries by identifying relevant passages from 45 million open-access papers. Works like DeepReviewer zhu2025deepreviewimprovingllmbasedpaper incorporate OpenScholar for novelty validation. However, our primary criticism of OpenScholar for novelty assessment is that it provides only generic comparisons rather than the granular analysis across methodology, problem formulation, evaluation approaches, and novelty claims that our task requires.
@@
Prior works evaluating generated peer reviews have adopted either quantitative evaluations, where they compare LLM-assigned scores (such as Overall Score, Soundness, etc.) against human-assigned scores on review forms, or qualitative evaluations using traditional metrics like BERTScore zhang2020bertscoreevaluatingtextgeneration, ROUGE lin-2004-rouge, and BLEU papineni-etal-2002-bleu, or more recent approaches like LLM-as-Judge zheng2023judging. We adopt the LLM-as-Judge approach for our evaluation. Notably, no prior work has specifically evaluated LLM performance on novelty assessment as a dedicated task, making our evaluation framework the first of its kind.
Dataset Scale: Our dataset comprises 182 papers and 352 reviews, which is comparable to or larger than datasets used in related peer review research: du-etal-2024-llms use 100 papers with 380 reviews, kennard-etal-2022-disapere label review data sourced for 188 papers, hua-etal-2019-argument work with 400 reviews, and chamoun-etal-2024-automated evaluate on 300 reviews. Novelty assessment requires careful manual annotation of scattered novelty discussions across reviews, making large-scale annotation resource-intensive. Following established practice in peer review analysis, we prioritize annotation quality over quantity, ensuring each example receives thorough annotation.
@@ 3.4 Stage 2: Related Work Discovery
We generate embeddings for all collected papers using SPECTER v2 Singh2022SciRepEvalAM on concatenated titles and abstracts. Papers are ranked by cosine similarity to the submission’s embedding to identify semantically similar work.
@@
To prioritize papers with conceptual rather than purely semantic similarity, we employ LLM-based reranking sun-etal-2023-chatgpt; Sun2023InstructionDM with prompts emphasizing methodological approaches, novelty claims, and problem statements. We select the top-K (k=20) papers for novelty assessment.
@@
For selected papers, we retrieve PDFs through a hierarchical search across Semantic Scholar, ACL Anthology, and arXiv. Retrieved papers are processed using MinerU wang2024mineruopensourcesolutionprecise; he2024opendatalab to extract introduction sections, with Nougat OCR blecher2023nougat as fallback for processing failures. We use these tools for OCRs here as they output more accurate OCRs and we will be using this paper content in the next stage.
@@ 3.5 Stage 3: Novelty Assessment
We use GPT-4.1 openai2024gpt41 for its improved instruction-following capabilities. This stage consists of four sequential steps.
@@
Processing retrieved papers as raw text creates context optimization challenges that degrade LLM performance. Recent research demonstrates that model performance consistently degrades with increasing input length, even when task complexity remains constant hong2025context. This occurs because either overwhelming models with unrelated information reduces accuracy zhu2025deepreviewimprovingllmbasedpaper; idahl-ahmadi-2025-openreviewer or insufficient context through heavy truncation limits understanding radensky2025scideatorhumanllmscientificidea.
@@ 4 Evaluation
OpenReviewer idahl-ahmadi-2025-openreviewer 42.4 $\pm$ 0.39 46.8 $\pm$ 0.71 6.3 $\pm$ 0.27 15.3 $\pm$ 0.40
DeepReviewer zhu2025deepreviewimprovingllmbasedpaper 50.6 $\pm$ 0.67 51.5 $\pm$ 1.24 21.7 $\pm$ 1.89 9.1 $\pm$ 0.00
@@ 4.3 Baseline Methods
OpenReviewer idahl-ahmadi-2025-openreviewer
@@
DeepReviewer zhu2025deepreviewimprovingllmbasedpaper
DeepReviewer is a multi-stage review framework that combines literature retrieval done with OpenScholar asai2024openscholarsynthesizingscientificliterature with evidence-based argumentation, powered by DeepReviewer-14B trained on structured review annotations. We extract novelty assessments using the same approach as OpenReviewer. Notably, DeepReviewer was trained on ICLR 2025 data, which includes our entire evaluation dataset.
@@ 5.1 Overall Performance
Our system significantly outperforms both AI baselines and the human-vs-human baseline. For Reasoning Alignment, we achieve scores 44.1 and 35.9 percentage points higher than OpenReviewer idahl-ahmadi-2025-openreviewer and DeepReviewer zhu2025deepreviewimprovingllmbasedpaper, and 21.4 points above the human baseline. For Conclusion Agreement, our system again leads all baselines, outperforming the nearest human baseline by approximately 13 percentage points (Tables 9, 10, 11).
@@ 5.3 Component Analysis
Table 3 shows the incremental contribution of each pipeline component. Our human-informed prompt design provides the largest gains (+40.7% reasoning, +46.8% conclusion), reflecting the importance of structured evaluation criteria derived from our human analysis. Structured extraction adds moderate improvements (+3.3% reasoning, +4.5% conclusion) but reduces overall computation costs and time, while landscape analysis contributes minimally (+3.2% reasoning, -0.7% conclusion). The major improvements come from the prompt design and this is an interesting find as it shows that with careful prompt design we are able to outperform the more complex method that underwent extensive training.
@@ 6 Conclusion
Our approach demonstrates that careful prompt design can achieve strong performance without requiring extensive model training. This is a strength of our method. While methods that attempt training zhu2025deepreviewimprovingllmbasedpaper; idahl-ahmadi-2025-openreviewer require substantial computational resources (e.g., 8× H100 80G GPUs for 23,500 steps at 256K context), whereas our prompt-based approach achieves strong performance while offering greater computational efficiency.
@@
Human Study: Our human evaluation is based on 100 pairwise comparisons with three expert annotators, comparable to related work in peer review analysis (Yuan_Liu_2022: 40 papers; chamoun-etal-2024-automated 100 examples; dycke-etal-2025-stricta: 20+ papers). While a larger sample would provide additional confidence, novelty assessment requires in-depth domain expert annotators with deep familiarity with the relevant literature, making extensive human evaluation resource-prohibitive.
System Surface-Level (%) Moderate (%) Deep (%)
OpenReviewer 67.4 31.3 1.2
DeepReviewer 43.4 56.6 0.0
Human vs. Human 22.3 66.2 11.5
Scideator 44.9 54.5 0.6
Ours 0.0 47.9 52.1
Table 4: Reasoning Depth Distribution (Percentages)
System None (%) Limited (%) Extensive (%)
OpenReviewer 39.9 53.1 7.0
DeepReviewer 24.7 75.3 0.0
Human vs. Human 19.6 65.2 15.2
Scideator 0.0 75.9 24.1
Ours 0.0 39.1 60.9
Table 5: Prior Work Engagement Distribution (Percentages)
Method Top-5 Top-10 Top-15 Top-20
Full Pipeline 1.000 1.000 1.000 1.000
Dense Only 0.612 $\pm$ 0.20 0.710 $\pm$ 0.18 0.748 $\pm$ 0.12 0.766 $\pm$ 0.10
KW Only 0.300 $\pm$ 0.27 0.317 $\pm$ 0.24 0.345 $\pm$ 0.22 0.349 $\pm$ 0.19
Table 6: Retrieval Pipeline Ablation Study
@@ Appendix D Understanding Human alignment patterns
LLMs are known to hallucinate references, which is precisely why our pipeline is specifically designed to be grounded through an extensive multi-step retrieval pipeline with multiple reranking stages. Each related paper alongside the submission is also used in our pipeline. To directly address whether our system’s citations "accurately support the conclusions or occasionally introduce factual inconsistencies," we conducted a systematic manual verification study.
We randomly sampled 50 factual claims from our novelty delta analysis outputs. Each claim was manually verified against the actual paper. Claims were categorized as ACCURATE, PARTIALLY-ACCURATE, INACCURATE, or CANNOT-VERIFY. As can be seen in table 8 that 96% of claims were accurate or partially accurate. The 14 partially accurate claims typically involved minor discrepancies such as incorrect author attribution (e.g., "Sun et al." instead of "Xu et al.") or year errors (e.g., 2024 instead of 2023), while the core method descriptions remained correct. Only 2 claims (4%) contained substantive factual errors. One misattributing a paper’s domain (text vs. image attacks) and one overstating a method’s historical significance. We will include this factuality verification study in the Appendix of the revised paper.

### v4

@@ Abstract
Novelty assessment is a central yet understudied aspect of peer review, particularly in high-volume fields like NLP where reviewer capacity is increasingly strained. We present a structured approach for automated novelty evaluation that models expert reviewer behavior through three stages: content extraction from submissions, retrieval and synthesis of related work, and structured comparison for evidence-based assessment. Our method is informed by analysis of human-written novelty reviews and captures key patterns such as independent claim verification and contextual reasoning. Evaluated on 182 ICLR 2025 submissions with human annotated reviewer novelty assessments, the approach achieves 86.5% alignment with human reasoning and 75.3% agreement on novelty conclusions—substantially outperforming existing LLM-based baselines. The method produces detailed, literature-aware analyses and improves consistency over ad hoc reviewer judgments. These results highlight the potential for structured LLM-assisted approaches to support more rigorous and transparent peer review without displacing human expertise. Data and code are made available.  [^1] https://ukplab.github.io/eacl2026-assessing-paper-novelty/
@@ 1 Introduction
The peer review system is collapsing under its own success. Two independent committees at NeurIPS 2021 disagreed on 23% of identical papers Beygelzimer et al. (2023)—a breakdown in consistency that signals deeper problems than mere capacity constraints. With manuscript submissions doubling every 15 years Larsen and von Ins (2010) and reviewers now handling 14 evaluations annually Díaz et al. (2024), the system’s 15 million annual reviewing hours Aczel et al. (2021) are producing increasingly unreliable outcomes.
@@
Among peer review tasks, novelty assessment stands out as one of the most problematic Ernst et al. (2021) Horbach and Halffman (2019). Novelty assessment requires reviewers to determine whether a submission makes sufficiently original contributions by identifying what specific advances it makes beyond existing work, evaluating whether these advances are significant enough to warrant publication, and verifying that the authors have accurately characterized their contributions relative to prior research. This knowledge-intensive process demands that reviewers maintain comprehensive awareness of related work across their field and can precisely distinguish between meaningful innovations and incremental modifications—a task that becomes exponentially more difficult as publication rates accelerate and research domains specialize. Overwhelmed reviewers often resort to superficial analyses, producing vague feedback like "not novel enough" without clear justification. The challenge compounds when reviewers encounter papers outside their specific expertise, leading to either overly conservative rejections or inadequate assessments that fail to catch incremental work Kuznetsov et al. (2024).
@@
Recent advances in large language models present an unprecedented opportunity to address these novelty assessment challenges at scale. These breakthrough technologies have revolutionized text processing and demonstrated remarkable performance across knowledge-intensive tasks Raiaan et al. (2024), with recent technical advancements expanding capabilities to specialized reasoning and efficient inference Li et al. (2024); Zhang et al. (2025).
@@
While recent LLM advances create this opportunity, no existing work specifically addresses novelty assessment as a dedicated task within the peer review process. Prior research incorporates novelty evaluation within idea generation pipelines Radensky et al. (2024); Lu et al. (2024); Li et al. (2025), generates peer reviews with novelty assessments occurring as a result of them existing in peer reviews from training data Idahl and Ahmadi (2025); D’Arcy et al. (2024), or adds novelty assessment steps to review synthesis pipelines for improvement Zhu et al. (2025). However, these approaches either operate on synthetic ideas rather than real research contributions or fail to evaluate novelty assessment capabilities in isolation. This represents a critical gap requiring specialized methodologies for peer review novelty assessment.
@@ 2 Related Work
Our work is positioned at the peer review stage of scientific research, where our system operates when a manuscript is submitted for evaluation. While previous works D’Arcy et al. (2024) Idahl and Ahmadi (2025) Zhu et al. (2025) Chitale et al. (2025) Chang et al. (2025) Nemecek et al. (2025) have developed end-to-end peer review generation pipelines that may implicitly include novelty assessment steps, we are the first to focus specifically on building a dedicated pipeline for novelty assessment and the first to systematically evaluate LLMs on this task. A related line of work operates at the ideation stage of research Radensky et al. (2024) Shahid et al. (2025) Li et al. (2025) Lu et al. (2024), developing pipelines for research idea generation that aim to improve novelty through feedback loops from a novelty assessor. In contrast, we operate at a more mature stage where ideas have been fully executed and comparative analyses are well-formulated. The evaluation in ideation-stage works focuses on synthetic ideas that are typically abstract and loosely defined, whereas we evaluate concrete, polished research contributions that have undergone the refinement process of execution and manuscript preparation.
@@
Our work employs an extensive related work discovery pipeline that collects papers cited within the submission and additionally retrieves related papers by querying with prompts generated by GPT-4.1. Papers are then ranked using an embedding-based method and reranked using RankGPT. We adapt this general approach from existing work Radensky et al. (2024)Shahid et al. (2025)Li et al. (2025) with modifications to ranking and filtering for our specific task. Similar retrieval-rank-rerank pipelines have been used for related work generation Agarwal et al. (2025). Another retrieval approach is OpenScholar Asai et al. (2024), which uses an LLM-RAG based approach to answer scientific queries by identifying relevant passages from 45 million open-access papers. Works like DeepReviewer Zhu et al. (2025) incorporate OpenScholar for novelty validation. However, our primary criticism of OpenScholar for novelty assessment is that it provides only generic comparisons rather than the granular analysis across methodology, problem formulation, evaluation approaches, and novelty claims that our task requires.
@@
Prior works evaluating generated peer reviews have adopted either quantitative evaluations, where they compare LLM-assigned scores (such as Overall Score, Soundness, etc.) against human-assigned scores on review forms, or qualitative evaluations using traditional metrics like BERTScore Zhang et al. (2020), ROUGE Lin (2004), and BLEU Papineni et al. (2002), or more recent approaches like LLM-as-Judge Zheng et al. (2023). We adopt the LLM-as-Judge approach for our evaluation. Notably, no prior work has specifically evaluated LLM performance on novelty assessment as a dedicated task, making our evaluation framework the first of its kind.
Dataset Scale: Our dataset comprises 182 papers and 352 reviews, which is comparable to or larger than datasets used in related peer review research: Du et al. (2024) use 100 papers with 380 reviews, Kennard et al. (2022) label review data sourced for 188 papers, Hua et al. (2019) work with 400 reviews, and Chamoun et al. (2024) evaluate on 300 reviews. Novelty assessment requires careful manual annotation of scattered novelty discussions across reviews, making large-scale annotation resource-intensive. Following established practice in peer review analysis, we prioritize annotation quality over quantity, ensuring each example receives thorough annotation.
@@ 3.4 Stage 2: Related Work Discovery
We generate embeddings for all collected papers using SPECTER v2 Singh et al. (2023) on concatenated titles and abstracts. Papers are ranked by cosine similarity to the submission’s embedding to identify semantically similar work.
@@
To prioritize papers with conceptual rather than purely semantic similarity, we employ LLM-based reranking Sun et al. (2023b, a) with prompts emphasizing methodological approaches, novelty claims, and problem statements. We select the top-K (k=20) papers for novelty assessment.
@@
For selected papers, we retrieve PDFs through a hierarchical search across Semantic Scholar, ACL Anthology, and arXiv. Retrieved papers are processed using MinerU Wang et al. (2024); He et al. (2024) to extract introduction sections, with Nougat OCR Blecher et al. (2024) as fallback for processing failures. We use these tools for OCRs here as they output more accurate OCRs and we will be using this paper content in the next stage.
@@ 3.5 Stage 3: Novelty Assessment
We use GPT-4.1 OpenAI (2024) for its improved instruction-following capabilities. This stage consists of four sequential steps.
@@
Processing retrieved papers as raw text creates context optimization challenges that degrade LLM performance. Recent research demonstrates that model performance consistently degrades with increasing input length, even when task complexity remains constant Hong et al. (2025). This occurs because either overwhelming models with unrelated information reduces accuracy Zhu et al. (2025); Idahl and Ahmadi (2025) or insufficient context through heavy truncation limits understanding Radensky et al. (2024).
@@ 4 Evaluation
OpenReviewer Idahl and Ahmadi (2025) 42.4 $\pm$ 0.39 46.8 $\pm$ 0.71 6.3 $\pm$ 0.27 15.3 $\pm$ 0.40
DeepReviewer Zhu et al. (2025) 50.6 $\pm$ 0.67 51.5 $\pm$ 1.24 21.7 $\pm$ 1.89 9.1 $\pm$ 0.00
@@ 4.3 Baseline Methods
OpenReviewer Idahl and Ahmadi (2025)
@@
DeepReviewer Zhu et al. (2025)
DeepReviewer is a multi-stage review framework that combines literature retrieval done with OpenScholar Asai et al. (2024) with evidence-based argumentation, powered by DeepReviewer-14B trained on structured review annotations. We extract novelty assessments using the same approach as OpenReviewer. Notably, DeepReviewer was trained on ICLR 2025 data, which includes our entire evaluation dataset.
@@ 5.1 Overall Performance
Our system significantly outperforms both AI baselines and the human-vs-human baseline. For Reasoning Alignment, we achieve scores 44.1 and 35.9 percentage points higher than OpenReviewer Idahl and Ahmadi (2025) and DeepReviewer Zhu et al. (2025), and 21.4 points above the human baseline. For Conclusion Agreement, our system again leads all baselines, outperforming the nearest human baseline by approximately 13 percentage points (Tables 9, 10, 11).
@@ 5.3 Component Analysis
Table 3 shows the incremental contribution of each pipeline component. Our human-informed prompt design provides the largest gains (+40.7% reasoning, +46.8% conclusion), reflecting the importance of structured evaluation criteria derived from our human analysis. Structured extraction adds moderate improvements (+3.3% reasoning, +4.5% conclusion) but reduces overall computation costs and time, while landscape analysis contributes minimally (+3.2% reasoning, -0.7% conclusion). The major improvements come from the prompt design and this is an interesting finding as it shows that with careful prompt design we are able to outperform the more complex method that underwent extensive training.
@@ 6 Conclusion
Our approach demonstrates that careful prompt design can achieve strong performance without requiring extensive model training. This is a strength of our method. While methods that attempt training Zhu et al. (2025); Idahl and Ahmadi (2025) require substantial computational resources (e.g., 8× H100 80G GPUs for 23,500 steps at 256K context), whereas our prompt-based approach achieves strong performance while offering greater computational efficiency.
@@
Human Study: Our human evaluation is based on 100 pairwise comparisons with three expert annotators, comparable to related work in peer review analysis (Yuan and Liu (2022): 40 papers; Chamoun et al. (2024) 100 examples; Dycke et al. (2025): 20+ papers). While a larger sample would provide additional confidence, novelty assessment requires in-depth domain expert annotators with deep familiarity with the relevant literature, making extensive human evaluation resource-prohibitive.
Acknowledgments
This work has been funded by the European Union (ERC, InterText, 101054961). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This work has also been co-funded by the LOEWE Distinguished Chair “Ubiquitous Knowledge Processing”, LOEWE initiative, Hesse, Germany (Grant Number: LOEWE/4a//519/05/00.002(0002)/81).
@@ Appendix D Understanding Human alignment patterns
LLMs are known to hallucinate references, which is precisely why our pipeline is specifically designed to be grounded through an extensive multi-step retrieval pipeline with multiple reranking stages. Each related paper alongside the submission is also used in our pipeline. To directly address whether our system’s citations accurately support the conclusions or occasionally introduce factual inconsistencies, we conducted a systematic manual verification study.
We randomly sampled 50 factual claims from our novelty delta analysis outputs. Each claim was manually verified against the actual paper. Claims were categorized as ACCURATE, PARTIALLY-ACCURATE, INACCURATE, or CANNOT-VERIFY. As can be seen in table 8 that 96% of claims were accurate or partially accurate. The 14 partially accurate claims typically involved minor discrepancies such as incorrect author attribution (e.g., "Sun et al." instead of "Xu et al.") or year errors (e.g., 2024 instead of 2023), while the core method descriptions remained correct. Only 2 claims (4%) contained substantive factual errors. One misattributing a paper’s domain (text vs. image attacks) and one overstating a method’s historical significance.t